A railway truck of 100 kg mass travelling with a velocity of 7 ms collides with a second truck of 200 kg mass and the two couple automatically and move off together. Calculate the velocity of the coupled trucks, and the energy lost to the system, if the second truck has a speed of 5 ms (al in the same direction, (b) in the opposite direction and ic) is initially stationary

1
Expert's answer
2022-02-20T15:45:20-0500

Given:

"m_1=100\:\rm kg"

"v_1=7\:\rm m/s"

"m_2=200\:\rm kg"

"v_2=5\:\rm m/s"


(a) The law of conservation of momentum

"m_1v_1+m_2v_2=(m_1+m_2)v""v=\frac{m_1v_1+m_2v_2}{m_1+m_2}=\frac{100*7+200*5}{100+200}=5.7\:\rm m/s"

(b) The law of conservation of momentum

"m_1v_1-m_2v_2=(m_1+m_2)v""v=\frac{m_1v_1-m_2v_2}{m_1+m_2}=\frac{100*7-200*5}{100+200}=-1.0\:\rm m/s"